# Distributional Bellman Operators over Mean Embeddings

### 09 Dec 2023 v1

@@ 7 Conclusion
We have proposed a framework for distributional reinforcement learning based on Bellman updates that take place entirely within the sketch domain. This has yielded new dynamic programming and temporal-difference learning algorithms as well as novel error propagation analysis, and we have provided further empirical analysis in the context of a suite of tabular MRPs, as well as demonstrating that the approach can be successfully applied at scale as a variant of the DQN architecture. We expect that there will be benefits from further exploration of algorithmic possibilities opened up by this framework, as well as potential consequences for value representations in the nervous system.
Acknowledgments
We thank Tim Genewein, Jiaxin Shi, Yayi Zou, Maneesh Sahani and Csaba Szepesvari for helpful discussions on the contractivity and generalisations of the sketch Bellman framework.
@@ B.1 Categorical imputation
B.2 Choices of feature function
  • Sigmoid: $\kappa(x)=\frac{1}{1+\exp(-x)}$;
  • Gaussian: $\kappa(x)=\exp(-x^{2}/2)$;
  • Hyperbolic tangent: $\kappa(x)=\tanh(x)$
  • Morbid: $\kappa(x)=\exp(-x^{2}/2)\cos(x)$
  • Triangular: $\kappa(x)=1-|x|$ for $|x|\leq 1$, zero otherwise.
  • Parabolic: $\kappa(x)=1-x^{2}$ for $|x|\leq 1$, zero otherwise.
  • Quartic: $\kappa(x)=(1-x^{2})^{2}$ for $|x|\leq 1$, zero otherwise.
In Appendix D.1, we provide further experimental results for a wide variety of feature maps from this family. In addition, we also consider the indicator feature used in Proposition 4.4.
@@ B.5 Mathematical properties of Bellman coefficients
B.6 Computational properties of Bellman coefficients
Rearranging, we obtain
@@ B.7 Comparison with other approaches to distributional RL
In this section, we provide additional comparisons against existing approaches to distributional RL, including their theoretical analysis. As distributional RL is a quickly evolving field, we focus our comparison on a few main classes of algorithms related to our work, which illustrate some key axes of variation within the field: (i) categorical approaches (Bellemare et al., 2017); (ii) quantile approaches (Dabney et al., 2018a; b; Yang et al., 2019); (iii) approaches related to maximum mean discrepancy (MMD; Gretton et al., 2012), such as Nguyen-Tang et al. (2021); Zhang et al. (2021); Sun et al. (2022); and (iv) sketch-based approaches (Sobel, 1982; Rowland et al., 2019).
@@
$\sum_{i=1}^{m}\frac{1}{m}\delta_{z_{i}}\,,$
in which the particle locations $(z_{i})_{i=1}^{m}$ are learnable. Work on sketches has instead focused on learning the values of particular statistical functionals of the return, rather than explicitly approximating return distributions. Rowland et al. (2019) also shows that standard categorical- and quantile-based algorithms can also be viewed through the lens of sketch-based distributional RL. The approach proposed in this paper sits firmly in the camp of sketch-based approaches, without ever representing approximated distributions directly. We highlight generative models of distributions as another prominent class of (non-parametric) representation (see, e.g., Doan et al. 2018; Freirich et al. 2019; Dabney et al. 2018a; Yang et al. 2019; Wu et al. 2023).
Algorithm types. Most prior algorithmic contributions to distributional reinforcement learning have focused on sample-based temporal-difference approaches, in which prediction parameters are iteratively and incrementally updated based on the gradient of a sampled approximation to a loss function. These approaches include the original C51 (Bellemare et al., 2017), QR-DQN (Dabney et al., 2018b), MMDRL (Nguyen-Tang et al., 2021), and EDRL (Rowland et al., 2019) algorithms. Dynamic programming algorithms, in which parameters are not updated incrementally via loss gradients, but instead according to the application of an implementable operator, have also been considered (see Rowland et al. (2018; 2023); Wu et al. (2023) for categorical dynamic programming, quantile dynamic programming, and fitted likelihood estimation, respectively). Marthe et al. (2023) discussed learning of certain risk-sensitive functionals in finite-horizon MDPs. In this paper, our algorithmic contributions include both DP and TD methods.
Losses and projections. One of the core axes of variation across distributional RL approaches is the loss used to define updates in incremental algorithms, and to define projections in dynamic programming. Categorical approaches use a projection in Cramér metric (Rowland et al., 2018) to define a target distribution for both dynamic programming and incremental versions of the algorithm; the incremental algorithm updates predictions via the gradient of a KL loss between the current and target distributions. Quantile-based approaches use either the quantile regression loss in incremental settings, or a Wasserstein-1 projection in dynamic programming (Dabney et al., 2018b). The MMDRL approach has been proposed only in the incremental setting, and considers sample-based gradients through an MMD loss, specifically taking the form
@@
for some choice of kernel $K$. In contrast, the Sketch-DP and Sketch-TD algorithms introduced in this paper define updates via matrix-vector products with the Bellman coefficients derived in Equation (5). With the connection to RKHS noted in Section 6, one can view the squared norm appearing in Equation (5) as the squared norm in the RKHS generated by the feature functions $\phi$. From this perspective, it is clear that both the Bellman coefficients that feature in Sketch-DP and Sketch-TD, and the gradients of the loss in Equation (11), both relate to the mathematical structure of RKHS.
@@
Theoretical analysis. The analysis of both dynamic programming and sample-based incremental algorithms requires analysis of the interaction between distributional updates, and approximations made to the distributions due to the choice of representation mentioned above.
Convergence analysis of dynamic programming algorithms has been obtained for several classes of distributional algorithms; see Rowland et al. (2018) for the case of categorical dynamic programming, Dabney et al. (2018b) for the case of a quantile dynamic programming algorithm, and Bellemare et al. (2023); Rowland et al. (2023) for later generalisations of this work. This analysis typically centres around (i) proving contractivity of the distributional Bellman operator $\mathcal{T}^{\pi}$ with respect to some metric d, and proving non-contractivity of the specific distributional projection used by the dynamic programming algorithm under this same metric. Notably, the metric d used in the analysis need not be the same as any metrics used in defining the algorithm; this is the case for quantile dynamic programming, for which Wasserstein-1 distance is used to define the algorithm, while Wasserstein-$\infty$ distance is used to analyse the algorithm (Dabney et al., 2018b; Bellemare et al., 2023; Rowland et al., 2023). Our proof technique, in particular, makes use of contraction of the distributional Bellman operator in Wasserstein distances, though such distances do not feature in the definition of the Sketch-DP/TD algorithms.
In general, there has been less work on the convergence analysis of sample-based incremental algorithms. Rowland et al. (2023) recently showed convergence of quantile temporal-difference learning, though the question of convergence for many other sample-based incremental distributional reinforcement learning algorithms is currently open. In particular, we note as a point of comparison to Nguyen-Tang et al. (2021), that although this work provides a contraction analysis of $\mathcal{T}^{\pi}$ (Theorem 2), and MMD approximation bounds for fixed target distributions (Theorem 3 and Proposition 2), these do not constitute a proof of convergence of the incremental algorithm described therein. The analysis of incremental algorithms is generally more mathematically involved than in the dynamic programming case, principally owing to the fact that rather than analysing the iterated application of a fixed operator, one needs to analyse the continuous dynamical system associated with incremental updates.
@@ C.2 Deep reinforcement learning implementation details
Baseline methods. We also tuned the number of atoms of the approximating distributions in the baseline methods. In particular, we found that C-51 performed the best compared to using more atoms; IQN did best when using 51 quantiles; and QR-DQN did best using 201 quantiles. Increasing the number of atoms in these methods lead to worse performance. We report the results of these best variants of the corresponding baseline methods in Figure 5.
@@ Appendix D Further experiments
D.1 Extended tabular results
@@
D.2 Extended results on Atari suite
@@
D.3 Temporal-difference learning with polynomial features
As noted in the main text, the sketch corresponding to the polynomial feature function $\phi(g)=(1,g,g^{2},\ldots g^{m})$ is Bellman closed, and the Bellman coefficients $B_{r}$ obtain zero regression error in Equation (5). In addition, there has been much prior work on dynamic programming (Sobel, 1982) for moments of the return, and temporal-difference learning specifically in the case of the first two moments (Tamar et al., 2013; 2016). However, such polynomial feature functions are difficult to use as the basis of learning high-dimensional feature embeddings. This stems from several factors, including that the typical scales of the coordinates of the feature function often vary over many orders of magnitude, making tuning of learning rates difficult, as well as the fact that polynomial features are non-local, making it more difficult to decode distributional information via an imputation strategy.
D.4 Comparison with statistical functional dynamic programming

### 04 Mar 2024 v3

@@ 7 Conclusion
We have proposed a framework for distributional RL based on Bellman updates that take place entirely within the sketch domain. This has yielded new dynamic programming and temporal-difference learning algorithms and a novel error propagation analysis. We have provided empirical validation on a suite of tabular MRPs and demonstrated that the approach can be successfully applied as a variant of the DQN. While convergence analysis for general sketches is an immediate future work, we expect that there will be benefits from further exploration of algorithmic possibilities opened up by this framework, and potential consequences for modelling value representations in the brain.
@@ B.1 Categorical imputation
B.2 Computational properties of Bellman coefficients
Rearranging, we obtain
@@ B.5 Mathematical properties of Bellman coefficients
B.6 Choices of feature function
  • Sigmoid: $\kappa(x)=\frac{1}{1+\exp(-x)}$;
  • Gaussian: $\kappa(x)=\exp(-x^{2}/2)$;
  • Parabolic: $\kappa(x)=1-x^{2}$ for $|x|\leq 1$, zero otherwise.
  • Hyperbolic tangent: $\kappa(x)=\tanh(x)$
In Appendix D.3, we provide further experimental results for a wide variety of feature maps from this family. In addition, we also consider the indicator feature used in Proposition 4.4 as well as sinusoidal features shown in Figure 2B.
@@ B.7 Comparison with other approaches to distributional RL
In this section, we provide additional comparisons against existing approaches to distributional RL. As distributional RL is a quickly evolving field, we focus our comparison on a few main classes of algorithms related to our work, which illustrate some key axes of variation within the field: (i) categorical approaches (Bellemare et al., 2017); (ii) quantile approaches (Dabney et al., 2018a, b; Yang et al., 2019); (iii) approaches related to maximum mean discrepancy (MMD; Gretton et al., 2012), such as Nguyen-Tang et al. (2021); Zhang et al. (2021); Sun et al. (2022); and (iv) sketch-based approaches (Sobel, 1982; Rowland et al., 2019).
@@
$\sum_{i=1}^{m}\frac{1}{m}\delta_{z_{i}}\,,$ (11)
in which the particle locations $(z_{i})_{i=1}^{m}$ are learnable. Work on sketches has instead focused on learning the values of particular statistical functionals of the return, rather than explicitly approximating return distributions.
Rowland et al. (2019) also shows that standard categorical- and quantile-based algorithms can also be viewed through the lens of sketch-based distributional RL. Other work in this vein includes Sobel (1982), who analysed the case of moments specifically, and Marthe et al. (2023), who extend the work of Rowland et al. (2019) to the undiscounted, finite-horizon case. The approach proposed in this paper sits firmly in the camp of sketch-based approaches, without ever representing approximated distributions directly. We highlight generative models of distributions as another prominent class of (non-parametric) representation (see, e.g., Doan et al. 2018; Freirich et al. 2019; Dabney et al. 2018a; Yang et al. 2019; Wu et al. 2023).
Algorithm types. Most prior algorithmic contributions to distributional reinforcement learning have focused on sample-based temporal-difference approaches, in which prediction parameters are iteratively and incrementally updated based on the gradient of a sampled approximation to a loss function. These approaches include the original C51 (Bellemare et al., 2017), QR-DQN (Dabney et al., 2018b), MMDRL (Nguyen-Tang et al., 2021), and EDRL (Rowland et al., 2019) algorithms. Dynamic programming algorithms, in which parameters are not updated incrementally via loss gradients, but instead according to the application of an implementable operator, have also been considered (see Rowland et al. (2018, 2023); Wu et al. (2023) for categorical dynamic programming, quantile dynamic programming, and fitted likelihood estimation, respectively). In this paper, our algorithmic contributions include both DP and TD methods.
Losses, projections, and convergence theory. One of the core axes of variation across distributional RL approaches is the loss used to define updates in incremental algorithms, and to define projections in dynamic programming. Categorical approaches use a projection in Cramér metric (Rowland et al., 2018) to define a target distribution for both dynamic programming and incremental versions of the algorithm; the incremental algorithm updates predictions via the gradient of a KL loss between the current and target distributions. Quantile-based approaches use either the quantile regression loss in incremental settings, or a Wasserstein-1 projection in dynamic programming (Dabney et al., 2018b).
The approach proposed in this paper works entirely with mean embeddings of probability distributions. Earlier approaches to sketched-based distributional RL, both in dynamic programming and incremental forms, have defined losses via imputation strategies, which compute updates by converting sketches into approximate distributions (Rowland et al., 2019; Bellemare et al., 2023).
Here, we contribute a novel perspective on the work of Nguyen-Tang et al. (2021), who propose a sample-based TD algorithm for updating particle locations (as in Equation (11)) by using an MMD loss, specifically taking the form
@@
for some choice of kernel $K$. Although not described in this manner by Nguyen-Tang et al. (2021), this can be seen as an incremental update on approximate mean embeddings for the RKHS $\mathcal{H}_{K}$ corresponding to the kernel $K$, amongst the class of mean embeddings of the form
$\sum_{i=1}^{m}\frac{1}{m}K(z_{i},\cdot)\in\mathcal{H}_{K}\,,$ (13)
where the particle locations $(z_{i})_{i=1}^{m}$ are optimised. We note that Nguyen-Tang et al. (2021) do not provide theoretical analysis for their algorithm. They provide contraction analysis of $\mathcal{T}^{\pi}$ (Theorem 2), and MMD approximation bounds for fixed target distributions (Theorem 3 and Proposition 2), but these do not constitute a proof of convergence of the incremental algorithm described therein. A key reason why the proposed algorithm may not yield clean convergence theory is that the space of mean embeddings described in Equation (13), where only the particles $(z_{i})_{i=1}^{m}$ can vary, is a non-convex subset of an infinite-dimensional RKHS. As such, tractable global optimisation of the objective may not be possible, along with the definition of a straightforward dynamic programming version of this approach. In other words, it is not straightforward to define a dynamic programming method to optimise the particle locations $(z_{i})_{i=1}^{m}$.
In contrast, the Sketch-DP and Sketch-TD algorithms introduced in this paper work with finite-dimensional RKHS, and define updates via matrix-vector products with the Bellman coefficients derived in Equation (5). This naturally yields tractable dynamic programming and temporal-difference learning algorithms, and also allows us to develop convergence theory, as described in Section 4.
@@
In general, convergence analysis of dynamic programming algorithms has been obtained for several classes of distributional algorithms beyond the theory described in this paper; see Rowland et al. (2018) for the case of categorical dynamic programming, Dabney et al. (2018b) for the case of a quantile dynamic programming algorithm, Bellemare et al. (2023); Rowland et al. (2023) for later generalisations of this work, and Wu et al. (2023) in the case of fitted likelihood evaluation. This analysis typically centres around (i) proving contractivity of the distributional Bellman operator $\mathcal{T}^{\pi}$ with respect to some metric d, and proving non-contractivity of the specific distributional projection used by the dynamic programming algorithm under this same metric. Notably, the metric d used in the analysis need not be the same as any metrics used in defining the algorithm; this is the case for quantile dynamic programming, for which Wasserstein-1 distance is used to define the algorithm, while Wasserstein-$\infty$ distance is used to analyse the algorithm (Dabney et al., 2018b; Bellemare et al., 2023; Rowland et al., 2023). Our proof technique in this paper, in particular, makes use of contraction of the distributional Bellman operator in Wasserstein distances, though such distances do not feature in the definition of the Sketch-DP/TD algorithms.
In general, there has been less work on the convergence analysis of sample-based incremental algorithms. Rowland et al. (2023) recently showed convergence of quantile temporal-difference learning, though the question of convergence for many other sample-based incremental distributional reinforcement learning algorithms is currently open. The analysis of incremental algorithms is generally more mathematically involved than in the dynamic programming case, principally owing to the fact that rather than analysing the iterated application of a fixed operator, one needs to analyse the continuous dynamical system associated with incremental updates.
@@ C.2 Deep reinforcement learning implementation details
Baseline methods. We also tuned the number of atoms of the approximating distributions in the baseline methods. In particular, we found that C51 performed the best compared to using more atoms; IQN did best when using 51 quantiles; and QR-DQN did best using 201 quantiles. Increasing the number of atoms in these methods lead to worse performance. We report the results of these best variants of the corresponding baseline methods in Figure 5.
@@ Appendix D Further experiments
D.1 Temporal-difference learning with polynomial features
As noted in the main text, the sketch corresponding to the polynomial feature function $\phi(g)=(1,g,g^{2},\ldots g^{m})$ is Bellman closed, and the Bellman coefficients $B_{r}$ obtain zero regression error in Equation (5). In addition, there has been much prior work on dynamic programming (Sobel, 1982) for moments of the return, and temporal-difference learning specifically in the case of the first two moments (Tamar et al., 2013, 2016). However, such polynomial feature functions are difficult to use as the basis of learning high-dimensional feature embeddings. This stems from several factors, including that the typical scales of the coordinates of the feature function often vary over many orders of magnitude, making tuning of learning rates difficult, as well as the fact that polynomial features are non-local, making it more difficult to decode distributional information via an imputation strategy.
D.2 Sketch DP example with sigmoidal features
ARefer to caption BRefer to caption
CRefer to caption
D                              ERefer to captionRefer to captionRefer to caption
Figure 8: An example run of Sketch-DP as in Figure 2 but using sigmoidal features as shown in B. We use $m=13$ sigmoidal features with anchors evenly spaced in $[-4.5,4.5]$ at locations indicated by the dotted lines. The regression Equation (5) is performed under a densely spaced grid over the white region $[-4,4]$.
To given an example of how to use the translation family described in Equation (8), we include another example of Sketch-DP execution in Figure 8. The results are largely similar to using sinusoidal features, except that the evolution is smoother and compared to using sinusoids. We also found that the first two principal components explained $>$97% variance, higher than using sinusoids. Please see Sections B.3 and B.6 for how to the heuristics on selecting the anchors and $\mu$.
D.3 Extended tabular results
@@
D.4 Comparison with statistical functional dynamic programming
D.5 Extended results on Atari suite
@@
D.6 Runtime comparison
Here, we report the mean (±s.d.) rate (per second) at which frames are processed during training in our Atari experiments, with each agent running on a single V100 GPU. These frame-processing times reflect the wallclock time associated with all aspects of the DQN training, including network forward passes for action selection, environment simulation, and periodic gradient updates. These statistics are averaged across all games and seeds.
The results are shown in Table 1. Note that, by default, C51 uses 51 atoms, QR-DQN uses 201 quantiles, and IQN uses 64 quantiles.
Number of features / atoms
Method 51 64 201 401
Sketch-DQN - - 1326 ± 107 1320 ± 110
Categorical-DQN (C51) 1309 ± 146 1306 ± 139 1293 ± 152 1291 ± 154
QR-DQN - - 1258 ± 107 1256 ± 106
IQN - 1120 ± 90 698 ± 41 400 ± 16
Table 1: Frame rate for selected methods. Higher is faster.
The Sketch-DQN method has the highest frame rate, and C51 and QR-DQN has slightly lower average frame rates. IQN with default 64 quantiles has a slightly lower average frame rate still, and is much lower when the number of quantiles is increased to match the number of predictions made by QR-DQN and Sketch-DQN. This is because the IQN architecture requires one forward pass through the MLP component of the network for each predicted quantile level. By contrast, the Sketch & QR architectures simply modify the original DQN architecture to produce multiple predictions of sketch values/quantiles from the final hidden layer of the network.
[…]